OpenStudy (anonymous):

Suppose R'(1)=16. a) If f(x)=R(5x), find f '(1/5) b) If g(x)=R(x+8), find g'(−7) c)If h(x) = R(x/7), find h'(7)

OpenStudy (gorv):

f ' (x)=R'(5x)*5 f ' (1/5)=R ' (5*1/5)*5=R'(1)*5 R'(1)=16 given f ' (1/5)=16*5=80

OpenStudy (gorv):

similarly do for other